OpenStudy (anonymous):

how do you solve this equation just show me the steps to the answer, (12)/(y)+(12)/(y+1)=1, the parentheses are just there to show you how the equation goes together.

OpenStudy (jamesj):

Multiply both sides by y(y+1) and have a look at what you get ...

OpenStudy (anonymous):

Show me I can't figure it out

OpenStudy (jamesj):

Multiplying both sides by y(y+1) 12y(y+1)/y + 12y(y+1)/(y+1) = y(y+1) So 12(y+1) + 12y = y^2 + y Following?

OpenStudy (anonymous):

Yes, Continue!

OpenStudy (jamesj):

So LHS = 12y + 12 + 12y = 24y + 12 Hence y^2 + y - (24y + 12) = 0 i.e., y^2 - 24y - 12 = 0 Yes? And now you can solve this quadratic equation

OpenStudy (anonymous):

How did you get 24y+12?

OpenStudy (jamesj):

The left hand side, the LHS = 12(y+ 1) + 12 = 12y + 12 + 12y = 24y + 12

OpenStudy (anonymous):

ooooo I understand now thanks!

OpenStudy (jamesj):

'good answer' if you don't mind

OpenStudy (jamesj):

What grade are you in?

OpenStudy (anonymous):

10th but I suck at math

OpenStudy (jamesj):

got it. Write out this example with all the steps by yourself until you get it. The start with a blank piece of paper and do it again. Repeat that until you can do it without looking. This way, you'll really learn this. Please also give me a 'good answer'

OpenStudy (anonymous):

Quick question first how did you get the 24y+12 out of the parentheses?

OpenStudy (jamesj):

12(y+ 1) + 12y = 12y + 12 + 12y = 24y + 12

OpenStudy (jamesj):

We had: 12(y+1) + 12y = y^2 + y The LHS = 12(y+ 1) + 12y = 12y + 12 + 12y = 24y + 12

OpenStudy (anonymous):

12/(y+1)+12/y = 1 Multiply both sides by y: (12 y)/(y+1)+12 = y Write the left hand side as a single fraction: (12 (2 y+1))/(y+1) = y Expand out terms of the left hand side: (24 y)/(y+1)+12/(y+1) = y Multiply both sides by y+1: 24 y+12 = y (y+1) Expand out terms on the right hand side: 24 y+12 = y^2+y Subtract (y^2+y) from both sides: -y^2+23 y+12 = 0 Solve the quadratic equation by completing the square: Divide both sides by -1: y^2-23 y-12 = 0 Add 12 to both sides: y^2-23 y = 12 Add 529/4 to both sides: y^2-23 y+529/4 = 577/4 Factor the left hand side: (y-23/2)^2 = 577/4 Take the square root of both sides: sqrt(y-23/2) = sqrt(577)/2 Eliminate the absolute value: y-23/2 = -sqrt(577)/2 or y-23/2 = sqrt(577)/2 Add 23/2 to both sides: y = 1/2 (23-sqrt(577)) or y-23/2 = sqrt(577)/2 Add 23/2 to both sides: y = 1/2 (23-sqrt(577)) or y = 1/2 (23+sqrt(577))

OpenStudy (anonymous):

Thanks @JamesJ